# Three-in-One: Fast and Accurate Transducer for Hybrid-Autoregressive ASR

### v1

@@ 5 Experiments
We evaluate HAINAN ASR in two languages: English, German. All experiments are conducted using the NeMo (Kuchaiev et al., 2019) toolkit, version 1.23.0. We first extract 80-dimensional filterbank features on audio, with 25 ms windows at 10 ms strides. All models use FastConformer encoders with the first three layers all performing 2X subsampling, therefore 8X subsampling in total. Both TDT and HAINAN models use durations [0-8], and the probability of masking the predictor is 0.5 for HAINAN models. BPE tokenizer (Sennrich et al., 2016; Kudo & Richardson, 2018) of size 1024 is used for text representation. For all experiments, we let models train for sufficient steps until validation performance degrades (no more than 150k training steps), and run model averaging on 5 best checkpoints to generate the final model for evaluation. For all datasets, we report ASR performance measured by WER and the inference speed of RNN-T, CTC, TDT and HAINAN models. For HAINAN models, we include results of AR, NAR and SAR modes, and use the notation “SAR-$n$” to indicate $n$ rounds of refinement runs).
@@ 5.1 English ASR
We train our English models on the combination of Librispeech (Panayotov et al., 2015), Mozilla Common Voice (Ardila et al., 2019), VoxPopuli (Wang et al., 2021), Fisher (Cieri et al., 2004), People’s Speech (Galvez et al., 2021), Wall Street Journal (Paul & Baker, 1992), National Speech Corpus (Koh et al., 2019), VCTK (Yamagishi et al., 2019), Multilingual Librispeech (Pratap et al., 2020), Europarl (Koehn, 2005) datasets, plus “anonymous” datasets. In total, our dataset consists of approximately 60,000 hours of speech.
The encoder uses FastConformer-XXL architecture with 42 layers of conformer blocks, each of which uses 8 heads of self-attention layers with model hidden dimension = 1024, totaling around 1.1b parameters. The convolutions in the conformers use kernel size = 9. The encoder was initialized with the public checkpoint https://hf.co/nvidia/parakeet-tdt-1.1b. For standard RNN-T, TDT and HAINAN models, their predictors consist of 2-layer LSTMs, with hidden dimension = 640. The joint network is a 2-layer feed-forward network with ReLU in between and with hidden dimension 1024. For HAINAN models, we also include two variations. (1) a model trained with stateless predictor network, where the predictor outputs the learned embedding of the last one word in the history context and, (2) a model train with non-zero durations [1-8].
We evaluate the trained models using the Huggingface ASR leaderboard (Srivastav et al., 2023), including the following datasets: AMI test (ami), Earnings22 test (e22), Gigaspeech test (giga), Librispeech test-clean and test-other (clean and other), Spgispeech test (spgi), Tedlium test (ted), and VoxPopuli test (vox). The results are shown in Table 1. Since there are multiple datasets, we direct readers’ attention to the average WER achieved by different models.
Table 1: HAINAN English ASR: accuracy (WER%) and time on Huggingface ASR leaderboard datasets. AR/NAR/SAR represent autoregressive/non-autoregressive and semi-autoregressive inference modes. The number after SAR represents the number of refinement runs. Decoding time (seconds) is measured on only librispeech-test-other using batch=1.
@@
HAINAN NAR 15.78 13.73 9.77 1.49 2.91 3.49 3.75 5.75 7.08 40
stateless SAR-1 15.77 13.72 9.72 1.49 2.84 3.48 3.74 5.71 7.06 42
@@
We can see that,
  • In AR modes, regardless of the nature of the predictor, HAINAN models achieve superior accuracy than RNN-T and TDT. Furthermore, while standard HAINAN models achieve parity in terms of efficiency with TDT, the HAINAN with stateless predictor achieves better speed than TDT model with no performance degradation.
  • In NAR modes, HAINAN models achieve consistently better accuracy than CTC, with similar inference speed. In particular, NAR stateless HAINAN achieves better accuracy than the AR TDT model, while running over 2X faster; it also outperforms RNN-T model, while being over 4X faster.
  • SAR further improves the accuracy of NAR inference with small computational overhead. In the case of standard HAINAN, each additional SAR refinement run adds around 3-4 seconds of processing for the whole dataset, while HAINAN with stateless predictor only requires 1-2 seconds per SAR run. This is significantly faster than fully autoregressive TDT runs which takes around 40 seconds more processing time in total. Also in both models, just one round of SAR makes the model catch up with TDT performance while running at least 2X faster.
@@ 5.2 German ASR
Our German models are trained on German Mozilla Common Voice, Multilingual Librispeech, and VoxPopuli datasets, totaling around 2000 hours. We use FastConformerLarge configuration of around 110m parameters  [^3] The complete hyper-parameter details can be found at https://github.com/NVIDIA/NeMo/blob/main/examples/asr/conf/fastconformer/fastconformer_rnnt_bpe.yaml, and evaluate the models on the German Multilingual Librispeech (MLS) and VoxPopuli (Vox) testsets. For training, all model encoders are initialized with public checkpoint https://hf.co/nvidia/stt_en_fastconformer_transducer_large. The results of those models are shown in Table 2.
We see similar trends compared to our English models, where HAINAN achieves similar accuracy and efficiency when running in AR modes compared to TDT models, and achieves significantly improved NAR accuracy compared to CTC models, while achieving the same inference speed. SAR further closes the accuracy gaps between AR and NAR, while adding small overheads, and still being around 4X faster than AR inference.
Table 2: German ASR: accuracy (WER%) and inference time (seconds) on Multilingual Librispeech and VoxPopuli. Decoding time is measured with batch=1.
@@ 6.1 NAR HAINAN’s Capability of Capturing Acoustic Ambiguity
In this subsection, we highlight an aspect of NAR HAINAN model that grants it greater modeling power compared CTC models: its ability to support more diverse model representation and multiple hypotheses across different time stamps, especially for potentially ambiguous speech. To illustrate this, we present a real-world example using an audio recording of the phrase “ice cream”. First, we decode the utterance using our English CTC model and extract the arg-max token per frame, resulting in:
@@
The CTC model output correctly represents the “ice cream” hypothesis, with each non-blank token appearing once, surrounded by blanks. This aligns with the well-documented peaky behavior of CTC (Zeyer et al., 2021). In contrast, the arg-max tokens emitted from our NAR HAINAN model are:
@@
Notably, the per-frame arg-max output not only encodes “ice cream” but also captures an alternative hypothesis: “i scream”. These hypotheses are dispersed across different time frames. For clarity, we have color-coded the tokens: orange for “ice cream”, blue for “i scream”, and green for shared tokens. It’s worth noting that these hypotheses are indeed acoustically very similar, and their different alignments to the audio demonstrate that the NAR HAINAN’s per-frame arg-max output provides a more accurate representation of the audio’s ambiguity.
6.2 Viterbi-based Decoding
To utilize the better representation learned by NAR HAINAN models as described in Section 6.1, we propose a Viterbi-based inference method. Given an audio sequence of length $T$, we view the output of the NAR HAINAN model as a directed acyclic graph (DAG) with $T+1$ nodes, while the first $T$ correspond to frames and the last acts as a special end-of-sentence (eos) token. For each frame, we only consider the arg-max token probability, which represents the weight of the first $T$ nodes (the eos node has a weight 0). The duration outputs at each frame correspond to transition weights into neighboring nodes. For example, in our models with max-duration 8, each node has at most 8 incoming and 8 outgoing connections connecting to its neighbors.
Given this setting, speech recognition on the input audio is equivalent to finding the best path of the DAG from the first node to the last, which can be efficiently solved by a simplified “Viterbi decoding” algorithm. After Viterbi decoding generates the tokens and time stamps, we can use parallelized refinement techniques to further improve the results, similar to what is done with SAR decoding. Appendix Section A.3 shows the details of the Viterbi algorithm. We report the initial NAR-Viterbi decoding results and results after one round of refinement with our English ASR models in Table 3. We also include AR results for easy reference. We see that for HAINAN stateless and HAINAN with all positive durations, NAR-Viterbi achieves on-par accuracy to AR decoding; and in all cases, Viterbi with additional refinement steps can match or outperform AR results. This corroborates our analysis in the previous section that NAR HAINAN models learn better representations and it benefits from diverse hypotheses in encoder outputs.
Table 3: HAINAN English ASR using Viterbi decoding: accuracy (WER%) on Huggingface ASR leaderboard datasets. NAR-Viterbi represents Viterbi decoding using NAR outputs. Viterbi-$n$ means NAR-Viterbi plus $n$ refinement runs.
model decoding ami e22 giga clean other spgi ted vox AVG
HAINAN AR 15.92 14.16 9.70 1.46 2.76 3.59 3.59 5.60 7.10
NAR-Viterbi 16.23 14.03 9.75 1.48 2.81 3.53 3.60 5.60 7.13
Viterbi-1 16.13 13.93 9.71 1.47 2.77 3.51 3.60 5.56 7.09
AR 15.69 13.64 9.68 1.49 2.82 3.47 3.74 5.72 7.03
HAINAN NAR-Viterbi 15.87 13.45 9.72 1.47 2.87 3.42 3.74 5.70 7.03
stateless Viterbi-1 15.86 13.42 9.69 1.47 2.82 3.41 3.74 5.69 7.01
AR 15.75 13.37 9.73 1.40 2.83 3.52 3.56 5.76 6.98
HAINAN NAR-Viterbi 15.88 13.38 9.73 1.41 2.86 3.48 3.48 5.65 6.98
no duration 0 Viterbi-1 15.79 13.21 9.68 1.39 2.83 3.46 3.50 5.60 6.93
@@ 6.3 Impact of Duration Configurations
In this subsection, we examine how duration configurations affect HAINAN model performance. We vary the max-durations in our German models and report their ASR performance. Table 4 presents the WER(%) in NAR mode for different max-duration settings. Our results reveal a clear trend: models with max-durations of 1 and 2 perform significantly worse, while those with max-durations of 4 and 8 yield similar, improved results. This observation aligns with our earlier analysis, suggesting that smaller maximal durations constrain the model’s ability to learn complex structures. Interestingly, max-duration 4 provides the best overall results, despite our choice of max-duration 8 for most experiments. We selected the latter to accommodate greater linguistic variability across datasets, as we did not extensively tune this parameter.
These findings corroborate our previous subsection’s conclusion: the ability to skip more frames enables the model to learn better representations of audio input. The improved performance with longer durations demonstrates the model’s enhanced capacity to capture and represent complex acoustic-linguistic patterns, contributing to more accurate speech recognition.
Table 4: German ASR results of NAR HAINAN models with different maximum supported durations.
@@ 7 Conclusion and Future Work
  • In AR mode, HAINAN achieves improved performance to TDT and RNN-T models.
@@ A.2 Comparison with Hybrid-TDT-CTC models
A.3 Viterbi decoding algorithm
The Algorithm for Viterbi decoding is shown in Algorithm 2. For simplicity, this Algorithm returns a “backtrack” structure that encodes the selected frames found by the best-path algorithm. The Algorithm’s key is between lines 7 and 14, where nested loops run over all (time, duration) pairs, and for each node, it finds the probability of the best path from the first node to that node using a dynamic programing algorithm. The Algorithm can be carried out efficiently with $O(tn)$ complexity, where $t$ is the input sequence length and $n$ is the number of durations. Because duration 0 is not allowed in NAR modes, we assume there’s no duration 0 in $N$.
Algorithm 2 Viterbi Decoding of HAINAN Models
1:  Input: acoustic input $x$, supported durations $N$
2:  enc = encoder(x)
3:  token-probs, duration-probs = joint(enc, dec=None) # of shapes [T, V] and [T, len(N)]
4:  best-prob-per-frame = max(token-probs, dim=-1) # of shape [T]
5:  best-prob = [0.0 for t in range(len(x) + 1)]
7:  for target in range(1, len(x) + 1) do
8:     for idx, n in enumerate(N) do
10:        alpha = best-prob[source]
11:        trans-prob = best-prob-per-frame[target] * duration-prob[source,idx]
12:        if alpha * trans-prob $>$ best-prob[target] then
13:           best-prob[target] = alpha * trans-prob
14:           backtrack[target] = source
15:  Return: backtrack
After the Algorithm, we can extract the time stamps corresponding to the best path from “backtrack”, and extract the arg-max token for those stamps as the NAR-Viterbi decoding output. Furthermore, we can run additional parallelized refinements using the initial output and the model’s predictor network, similar to what is done with SAR decoding, to further improve the model output.

### v3

@@ 5 Experiments
We evaluate HAINAN ASR in two languages: English, German. All experiments are conducted using the NeMo (Kuchaiev et al., 2019) toolkit, version 1.23.0. We first extract 80-dimensional filterbank features on audio, with 25 ms windows at 10 ms strides. All models use FastConformer encoders with the first three layers all performing 2X subsampling, therefore 8X subsampling in total. Both TDT and HAINAN models use durations {0, 1, 2, …, 7, 8}. BPE tokenizer (Sennrich et al., 2016; Kudo & Richardson, 2018) of size 1024 is used for text representation. For all experiments, we let models train for sufficient steps until validation performance degrades (no more than 150k training steps), and run model averaging on 5 best checkpoints to generate the final model for evaluation. For all datasets, we report ASR performance measured by WER and the inference speed of RNN-T, CTC, TDT and HAINAN models. For HAINAN models, we include results of AR, NAR and SAR modes, and use the notation “SAR-$n$” to indicate $n$ rounds of refinement runs). We include Transducer models with stateless predictors with 1-token context denoted as “stateless”, and use “/0” to denote TDT/HAINAN models that use all positive durations {1, 2, …, 7, 8}.
@@ 5.1 English ASR
We train our English models on the combination of Librispeech (Panayotov et al., 2015), Mozilla Common Voice (Ardila et al., 2019), VoxPopuli (Wang et al., 2021), Fisher (Cieri et al., 2004), People’s Speech (Galvez et al., 2021), Wall Street Journal (Paul & Baker, 1992), National Speech Corpus (Koh et al., 2019), VCTK (Yamagishi et al., 2019), Multilingual Librispeech (Pratap et al., 2020), Europarl (Koehn, 2005) datasets, plus Suno AI datasets. In total, our dataset consists of approximately 60,000 hours of speech. The encoder uses FastConformer-XXL architecture with 42 layers of conformer blocks, each of which uses 8 heads of self-attention layers with model hidden dimension = 1024, totaling around 1.1b parameters. The convolutions in the conformers use kernel size = 9. The encoder was initialized with the public checkpoint https://hf.co/nvidia/parakeet-tdt-1.1b. For standard RNN-T, TDT and HAINAN models, their predictors consist of 2-layer LSTMs, with hidden dimension = 640. The joint network is a 2-layer feed-forward network with ReLU in between and with hidden dimension 1024.
We evaluate the models using the Huggingface ASR leaderboard (Srivastav et al., 2023), including the following datasets: AMI test (ami), Earnings22 test (e22), Gigaspeech test (giga), Librispeech test-clean and test-other (clean and other), Spgispeech test (spgi), Tedlium test (ted), and VoxPopuli test (vox). The results are shown in Table 1. Since there are multiple datasets, we direct readers’ attention to the average WER achieved by different models. From the Table we can see:
Table 1: HAINAN English ASR: accuracy (WER%) and wall time on Huggingface ASR leaderboard datasets. AR/NAR/SAR represent autoregressive/non-autoregressive and semi-autoregressive. The number after SAR represents the number of refinement runs. Decoding time (seconds) is measured on only librispeech-test-other using batch=1 and beam=1, running on 2 A6000 GPUs. Viterbi time is not included since it’s not comparable due to the different nature of the search algorithm. All models have around 1.1 billion parameters differing by at most 600k in actual number.
@@
NAR 15.78 13.73 9.77 1.49 2.91 3.49 3.75 5.75 7.08 40
stateless +Viterbi 15.87 13.45 9.72 1.47 2.87 3.42 3.74 5.70 7.03 -
HAINAN SAR-1 15.77 13.72 9.72 1.49 2.84 3.48 3.74 5.71 7.06 42
+Viterbi 15.86 13.42 9.69 1.47 2.82 3.41 3.74 5.69 7.01 -
@@
  • Regardless of the nature of the predictor, AR HAINAN models achieve superior accuracy than RNN-T and TDT counterparts, with similar efficiency.
  • NAR-HAINAN achieves consistently better accuracy than CTC, with similar inference speed. In particular, NAR stateless HAINAN achieves better accuracy than the AR TDT model, while running over 2X faster; it also outperforms RNN-T model, while being over 4X faster.
  • SAR further improves the accuracy of NAR inference with small computational overhead. With standard HAINAN, each additional SAR refinement run adds around 3-4 seconds of processing for the whole dataset, while stateless HAINAN only requires 1-2 seconds per SAR run. This is significantly faster than fully AR TDT runs which takes around 40 seconds more processing time in total. Also in both models, just one round of SAR makes the model catch up with TDT performance while running at least 2X faster.
  • Viterbi decoding can further improve model accuracy. We present an analysis in subsection 6.2 that can explain this phenomenon.
@@ 5.2 German ASR
Our German models are trained on German Mozilla Common Voice, Multilingual Librispeech, and VoxPopuli datasets, totaling around 2000 hours. We use FastConformerLarge configuration of around 110m parameters  [^3] The hyper-parameters can be found at https://github.com/NVIDIA/NeMo/blob/main/examples/asr/conf/fastconformer/fast-conformer_transducer_bpe.yaml, and evaluate the models on the German Multilingual Librispeech (MLS) and VoxPopuli (Vox) testsets. For training, all model encoders are initialized with public checkpoint https://hf.co/nvidia/stt_en_fastconformer_transducer_large. The results of those models are shown in Table 2.
We see similar trends compared to our English models, where HAINAN achieves similar AR accuracy and efficiency with TDT models, and significantly improved NAR accuracy compared to CTC models, with similar inference speed. SAR further closes the accuracy gaps between AR and NAR, while adding small overheads, and still being around 4X faster than AR inference.
Table 2: German ASR: accuracy (WER%) and inference wall time (seconds) on Multilingual Librispeech and VoxPopuli. Decoding time is measured with batch=1 and beam=1.
@@ 6 Analysis and Discussions
6.1 Limitations of SAR methods
The key of SAR method is to reuse the time stamps generated in the initial NAR inference, and only regenerates the token outputs for those frames. This puts limitations on the types of errors that can be corrected by the method. We provide some analysis here.
  • From the subword level, SAR can only fix substitution errors by replacing a wrong subword with the correct one, and insertion errors by replacing the inserted wrong subword with a blank. It can’t fix deletions where the hypothesis is shorter than the reference.
  • From the word level, SAR can correct deletion errors. Note, although SAR can’t add more subwords to the original output, it can replace a non-word-beginning subword with a word-beginning subword, and thus increase the number of words in the SAR output. E.g. the SAR can change “forty” to “for tea”, by replacing subwords “_for ty” with “_for _tea”, where “_” means the begin-of-word symbol.
Table 3: Breakdown of Error types (%) of German HAINAN Models on Voxpopuli Dataset.
inference sub del ins total
NAR 5.19 2.64 1.80 9.63
SAR-1 5.00 2.85 1.62 9.47
SAR-2 4.96 2.87 1.61 9.44
We report detailed breakdown of error types in our German HAINAN model on Voxpopuli comparing NAR and SAR in Table 3. We see that SAR models primarily reduce substitution and insertion errors; in fact, the total amount of deletion errors increased with SAR models. Even though the overall deletion error numbers go up, we do observe in certain utterances, SAR can correct deletion errors from NAR outputs. We provide some examples in Appendix subsection A.3.
@@ 6.2 NAR HAINAN’s Capability of Capturing Acoustic Ambiguity
In this subsection, we highlight an aspect of NAR HAINAN model that grants it greater modeling power compared CTC models: its ability to support more diverse model representation and multiple hypotheses across different time stamps, especially for potentially ambiguous speech. To illustrate this, we present a real-world example using an audio recording of the phrase “ice cream”. First, we run our English CTC model on the utterance to extract the arg-max token per frame, resulting in:
@@
The CTC model output correctly represents the “ice cream” hypothesis, with each non-blank token appearing once, surrounded by blanks. This aligns with the well-documented peaky behavior of CTC (Zeyer et al., 2021). In contrast, the arg-max tokens emitted from NAR HAINAN model are:
@@
Notably, the per-frame arg-max output not only encodes “ice cream” but also captures an alternative hypothesis: “i scream”. These hypotheses are dispersed across different time frames. For clarity, we have color-coded the tokens: orange for “ice cream”, blue for “i scream”, and green for shared tokens. It’s worth noting that these hypotheses are acoustically similar, and their different alignments to the audio demonstrate that the NAR HAINAN’s per-frame arg-max output provides a more accurate representation of the audio input. We believe this is the reason why the Viterbi algorithm can achieve improved accuracy.
@@ 6.3 Impact of Duration Configurations
In this subsection, we examine how duration configurations affect HAINAN model performance. We vary the max-durations in our German models and report their ASR performance in Table 4. Our results reveal a clear trend: models with max-durations of 1 and 2 perform significantly worse, while 4 and 8 yield improved results. This observation aligns with our earlier analysis, suggesting that shorter maximal durations constrain the model’s ability to learn complex structures, and the max duration needs to be sufficiently large to represent the real distribution. Interestingly, max-duration 4 provides the best overall results, despite our choice of max-duration 8 for most experiments. We selected the latter to accommodate greater linguistic variability across datasets, as we did not extensively tune this parameter.
Table 4: German ASR results of NAR HAINAN models with different maximum durations.
@@ 7 Conclusion and Future Work
  • In AR mode, HAINAN achieves on-par performance with TDT and RNN-T models.
@@
  • We also propose a simplified Viterbi decoding method to further improve model accuracy.
@@ A.2 Comparison with Hybrid-TDT-CTC models
A.3 Examples of deletion errors corrected by SAR methods
Although relatively rare, SAR methods can correct deletion errors of NAR inference results. Here are some examples in German and English that we observe in our runs. We use green to highlight words that were missing in NAR hypotheses but included by SAR.
German:
Table 6: Example of German ASR where SAR reduces deletion errors from NAR output
reference die meisten vorfälle passieren bei der implementierung bedauerlicherweise im krankenhaus häufig auch in der anwendung
NAR diesfälle passieren bei der implementierung bedauerlicherweise im krankenhaus häufig auch in der anwendung
SAR die handfälle passieren bei der implementierung bedauerlicherweise im krankenhaus häufig auch in der anwendung
reference ich ersuche den kommissar hier sofortmaßnahmen zu ergreifen
NAR ichsuche den kommissar hier sofort maßnahmen zu ergreifen
SAR ich versuche den kommissar hier sofortmaßnahmen zu ergreifen
English:
Table 7: Example of English ASR where SAR reduces deletion errors from NAR output
reference then one of them says kind of soft and gentle
NAR then one of them says kinda soft and gentle
SAR then one of them says kind of soft and gentle
reference say mester gurr sir which thankful i am to you for speaking so but you don’t really think as he has come to harm
NAR say mister gurrsir which thankful i am for you for speaking so but you don’t really think as he has come to harm
SAR say mister gurr sir which thankful i am for you for speaking so but you don’t really think as he has come to harm
We notice that, in all those examples of corrected deletion errors, the SAR method replaces a non-word-beginning subword with a word-beginning subword, e.g. “_kind a” is replaced with “_kind _of”. This aligns with our analysis in subsection 6.1, since this is the only way to increase the number of words while keeping the number of subwords fixed.
A.4 Reason of Zero-duration Mitigation with HAINAN
In subsection 6.4, we show empirically that duration 0 is greatly mitigated by HAINAN model. In this subsection, we provide some analysis of why this is the case. In our description, we use $E_{t}$ and $D_{u}$ to denote encoder/predictor outputs.
  1. 1.
    By performing stochastic masking of predictor output during training, the model is learning to make the output distribution of joint$(E_{t},D_{u})$ similar to joint$(E_{t},D_{u}*0)$.
  2. 2.
    If duration 0 is predicted from joint$(E_{t},D_{u})$, then the next step of processing would be joint$(E_{t},D_{u+1})$, where $t$ remains the same but $u$ gets incremented by 1.
  3. 3.
    From 1, we know that model is learning to make joint$(E_{t},D_{u})$ similar to joint$(E_{t},D_{u}*0)$; also it’s learning to make joint$(E_{t},D_{u+1})$ similar to joint$(E_{t},D_{u+1}*0)$. Consequently, the model would make joint$(E_{t},D_{u})$ similar to joint$(E_{t},D_{u+1})$, since joint$(E_{t},D_{u}*0)$ is the same as joint$(E_{t},D_{u+1}*0)$, both of which has predictor masked out.
  4. 4.
    However, we know that joint$(E_{t},D_{u})$ and joint$(E_{t},D_{u+1})$ should predict different tokens since the text histories are different. This causes a contradiction.
Therefore, HAINAN models, by adopting stochastic predictor masking, would naturally suppress the prediction of duration 0 in its processing.